# Enabling Delayed-Full Charging Through Transformer-Based Real-Time-to-Departure Modeling for EV Battery Longevity

### v1

@@ Introduction
In fact, the practical efficacy of the DFC algorithm relies accurate prediction of EV departure time. Inaccurate predictions introduce two critical risks: early charging reduces DFC ’s benefits by prolonging high SOC exposure, whereas late charging can leave insufficient energy at departure, causing range anxiety (Figure  1). Consequently, departure time emerges as a key control parameter for charging scheduling within the DFC framework. Existing methods primarily rely on historical statistics of usage patterns, offering limited adaptability to dynamic user contexts (frendo2021smart; lindroth2024online). These approaches often overlook real-time causal signals of departure, leading to poor performance and interpretability. For robust deployment of DFC, predictive models must provide temporal precision while adapting to contextual signals that reflect pre-departure behavior.
Pre-departure behavior is shaped by diverse contextual cues, including user behavioral signals such as physical activity, smartphone engagement, and social interactions, as well as environmental factors like ambient light and sound. Smartphone passive sensing provides a scalable and unobtrusive way to capture these signals continuously, enabling a digital phenotyping approach aligned with the quantified-self paradigm (torous2016new). Prior work has shown that machine learning classifiers trained on sliding-window contextual features can improve first-daily departure prediction from home. Combined with DBSCAN theses models outperform historical-pattern baselines  (lee2024extending). Yet, framing departure prediction as classification leads to severe label imbalance, which becomes more pronounced at finer time resolutions for real-time inference.
Building on recent advances in deep survival learning across domains such as healthcare, churn prediction, and mobility, we reformulate departure time prediction as a time-to-event (TTE) problem, applying survival analysis principles to estimate the probability distribution of departure. This reframing offers two key advantages: (1) it mitigates label imbalance by representing each day as a single observation instead of multiple binary windows, and (2) it supports real-time inference where the event time is unknown until departure, enabling probabilistic predictions. Unlike classification, TTE models predict survival functions, providing richer temporal modeling of departure likelihood. To this end, we propose a Time-to-Departure (TTD) modeling approach based on a Transformer architecture, leveraging its ability to capture temporal dependencies and achieve efficient parallel processing through multi-head attention. The model updates survival probabilities over discretized time grids using real-time contextual signals from passive sensing. This enables timely departure prediction and providing a robust foundation for successful deployment of DFC in real-world environments. Our contributions are threefold:
@@
This discrete-time formulation motivates estimating the sequence of hazard complements $\{q(t\mid\mathbf{X})\}_{t=0}^{T_{\max}}$ in a context-aware manner for each time step.
@@
Our framework predicts $\hat{q}(t\mid\mathbf{X})$ for all intervals in parallel using a Transformer-based architecture, enabling efficient survival estimation with real-time updates. Through self-attention, the model captures both regular daily routine and dynamically changing contextual patterns, supporting timely and adaptive departure predictions (hu2021transformer).
Objective. Given contextual observations up to time $t$, $\mathcal{X}_{0:t}$, the goal is to estimate the sequence of hazard complements $\{q(\tau\mid\mathbf{X})\}_{\tau=0}^{T_{\max}}$, which determines survival probabilities and departure likelihoods across the day. Formally, the model produces
$\hat{q}_{0:T_{\max}}=f_{\theta}(\mathcal{X}_{0:t}),$
where $f_{\theta}$ is the Transformer-based architecture parameterized by $\theta$. During training, supervision is applied only up to the observed departure time $t^{*}$, modeling survival for $t<t^{*}$ and failure at $t=t^{*}$. During inference, the model processes partial sequences in real time, updating survival estimates and triggering departure detection once the predicted survival probability falls below a decision threshold $\tau$. Figure 2 illustrates this real-time update process.
@@
Output Layer. For each interval $t$, the corresponding Transformer output $\mathbf{z}_{t}\in\mathbb{R}^{d_{\text{model}}}$ is mapped through a feed-forward network and sigmoid activation:
$\hat{q}(t|\mathbf{X})=\sigma(W_{o}\mathbf{z}_{t}+b_{o}),$
where $\hat{q}(t|\mathbf{X})\in(0,1)$ represents the predicted complement of the hazard probability. These outputs enable real-time estimation of survival probabilities as contextual observations evolve during the day.
Figure 2: Real-time survival modeling with streaming contextual inputs. During training, the model estimates survival probabilities up to the observed departure event for uncensored sequences. At inference, survival estimates are updated incrementally as new tokens arrive, triggering a departure once the probability falls below a threshold ($\tau$).
@@
We optimize a discrete-time ordinal regression objective inspired by (hu2021transformer). Let the training set contain $N$ user-day sequences indexed by $i\in\{1,\dots,N\}$. For each sequence $i$, let $\mathbf{X}_{i}$ denote the full contextual feature sequence and let $t^{*}_{i}\in\{0,\dots,T_{\max}\}$ denote the observed (uncensored) departure time. The model outputs hazard complements $\hat{q}(t\mid\mathbf{X}_{i})$ for $t\in[0,T_{\max}]$, and the survival probability up to time $t$ is:
$\hat{S}(t\mid\mathbf{X}_{i})=\prod_{\tau=0}^{t}\hat{q}(\tau\mid\mathbf{X}_{i}).$
This survival estimate forms the basis for the ordinal regression loss used to supervise the model up to the observed event time $t_{i}^{*}$.
Ordinal Regression Loss. Since all training sequences contain an observed departure event, we optimize the negative log-likelihood for uncensored samples:
$\mathcal{L}_{\text{ord},i}=-\sum_{t=0}^{\,t^{*}_{i}-1}\log\hat{q}(t\mid\mathbf{X}_{i})-\log\bigl(1-\hat{q}(t^{*}_{i}\mid\mathbf{X}_{i})\bigr),$
where the first term encourages survival up to $t^{*}_{i}$ and the second enforces failure at $t^{*}_{i}$. Intervals after $t^{*}_{i}$ are masked because post-event predictions are not relevant in TTD modeling. Since non-departure days represent true non-events rather than right-censored samples, training is performed solely on uncensored sequences.
Soft Supervision with Gaussian Smoothing. Short adjacent intervals (i.e., 5 minutes) may share similar context, making strict supervision sensitive to minor timing shifts near the event. To mitigate this, we apply a normalized Gaussian weighting kernel centered at $t^{*}_{i}$, restricted to pre-event intervals, to smooth the loss around the departure time. This Gaussian-Smoothed Supervision (GSS) enhances robustness to small label uncertainties and contextual noise immediately preceding the event:
$w_{i}(t)=\frac{\exp\bigl[-(t-t^{*}_{i})^{2}/(2\sigma^{2})\bigr]}{\sum_{\tau\leq t^{*}_{i}}\exp\bigl[-(\tau-t^{*}_{i})^{2}/(2\sigma^{2})\bigr]},\qquad t\leq t^{*}_{i},$
where $\sigma$ controls the smoothing range. The smoothed ordinal regression loss is
$\mathcal{L}_{\text{smooth},i}=-\sum_{t=0}^{t^{*}_{i}}w_{i}(t)\Big[$ $\mathbf{1}_{t<t^{*}_{i}}\log\hat{q}(t\mid\mathbf{X}_{i})$
${}+\mathbf{1}_{t=t^{*}_{i}}\log\bigl(1-\hat{q}(t\mid\mathbf{X}_{i})\bigr)\Big].$
Event and Weekend Weights. To enhance predictive accuracy, we incorporate two weighting factors: (1) the event weight $\omega_{e}$, which amplifies the failure term at $t^{*}_{i}$ and scales the corresponding gradients; and (2) the weekend weight $\omega_{w}$, which adjusts sequence-level contributions to account for irregular weekend patterns and stabilize the overall loss.
@@
$\text{where }\ell_{i,t}=\begin{cases}-\log\hat{q}(t\mid\mathbf{X}_{i}),&t<t^{*}_{i},\\[4.0pt] -\omega_{e}\log\bigl(1-\hat{q}(t^{*}_{i}\mid\mathbf{X}_{i})\bigr),&t=t^{*}_{i},\\[4.0pt] 0,&t>t^{*}_{i},\end{cases}$
@@
Figure 3: A diagram of TTD model architecture. Contextual features, absolute time indicators, and day-of-week embeddings are fused using alpha-fusion mechanism, with absolute time scaled by a learnable parameter $t$ to regulate its influence. The fused embeddings, combined with positional encoding, are processed by a multi-layer Transformer encoder to capture long-range temporal dependencies. The output layer predicts hazard complements $\hat{q}(t\mid\mathbf{X})$ across all intervals in parallel during training, and supports real-time inference by updating predictions as new tokens arrive.
@@
Evaluation Metric. Performance is measured using mean absolute error (MAE) in hours, which provides an intuitive measure of temporal deviation. The predicted departure time is taken as the earliest time index at which the estimated survival probability falls below a decision threshold. We vary this threshold from 0.05 to 0.20 in increments of 0.05 to assess its impact on accuracy, with 0.10 selected as the optimal value through hyperparameter tuning.
@@
Distribution Analysis. Figure 4 presents kernel density estimates (KDE) of predicted versus ground-truth departure times. Historical baselines exhibit high concentration around global averages, particularly in morning and afternoon periods, failing to capture contextual routines relevant to departure. In contrast, our model’s distribution closely tracks the ground-truth curve, demonstrating temporal adaptability despite minor deviations. A slight bias toward earlier predictions is observed, primarily due to the decision threshold. This suggests that adjusting the threshold improves practicality of DFC by providing a buffer time before departure.
@@
Figure 5: The impact of hyperparameters: event weight $\omega_{e}$, weekend weight $\omega_{w}$, and survival probability threshold $\tau$ on the departure time prediction accuracy.
@@
To evaluate the contribution of key design components, we conduct an ablation study along three dimensions: (1) input features, (2) architectural and regularization strategies, and (3) loss design. Table 2 presents MAE (hours) for all days, weekdays, and weekends (holidays) across these variants.
Feature Contributions. Removing contextual features (w/o context) leads to the largest performance drop (4.47 hours), underscoring the critical role of behavioral and environmental indicators in capturing pre-departure routines. Removing absolute time features (w/o time) results in moderate degradation (3.01 hours), indicating that temporal position cues are important for modeling heterogeneous patterns across days and individuals. Excluding day-of-week features (w/o DoW) slightly worsens performance (2.64 hours), highlighting their usefulness in capturing weekly rhythms.
@@
Loss Design and Synergistic Regularization. Removing both time-scale and GSS (w/o t-scale & GSS) increases MAE to 2.36 hours, confirming their combined benefit. Time-scale improves temporal calibration across discretized grids, while GSS provides soft supervision that reduces over-penalization of near-boundary timing shifts.
@@
We examine the sensitivity of our model to three key hyperparameters: event weight $\omega_{e}$, weekend weight $\omega_{w}$, and the survival probability threshold $\tau$. Figure 5 presents the MAE heatmaps across different combinations.
Effect of Event Weight. Each subplot corresponds to a fixed value of $\omega_{e}$. Medium values (1.0–1.5) yield the best MAE overall, especially when paired with moderate $\omega_{w}$. Small values weaken the event signal, while large values overemphasize the event point and harm calibration. Notably, the model’s stable performance at $\omega_{e}=1$ suggests that balancing survival and failure signals promotes robust training and stable outcomes. The best results are achieved when $\omega_{e}=1.5$ is combined with $\omega_{w}=1.5$ and a small $\tau$.
Effect of Weekend Weight. Small $\omega_{w}$ weakens weekend learning, causing the model to fall back on regular weekday patterns and making weekend cues less reliable. With large $\omega_{e}$ and $\tau$, these weak early cues combined with a softened threshold lead to unstable early predictions. With small $\omega_{e}$ and $\tau$, the event signal is underemphasized and the strict threshold delays detection, producing late predictions. Thus, small $\omega_{w}$ consistently harms weekend performance, while $\omega_{e}$ and $\tau$ determine whether errors occur early or late.
Effect of Survival Probability Threshold. $\tau$ shows a consistent MAE pattern centered around $\omega_{e}=1$, where survival and failure signals are balanced. When $\omega_{e}=1.0\text{–}2.0$, smaller $\tau$ improves accuracy since the model learns sharper survival drops near the event. When $\omega_{e}=0.5$, smaller $\tau$ delays detection and worsens performance. In short, strict $\tau$ works well when the event signal is emphasized.
Interactions and Practical Settings. Overall, moderate $\omega_{e}$ and $\omega_{w}$ (around 1.5) combined with a small $\tau$ yield the most reliable performance. Small $\omega_{w}$ consistently harms weekend predictions, while extreme $\omega_{e}$ or $\tau$ shifts the model toward unstable early or late detections. Using moderate weights and a strict threshold avoids these issues and provides accurate, stable real-time departure time predictions.
@@
We evaluate user-specific adaptation through fine-tuning. All global model parameters remain frozen, and only the last Transformer layer and output layer are updated using user-specific data. Personalization yields a modest overall improvement (MAE: 2.20 → 2.13 hours), with larger gains on weekends (2.07 → 1.85 hours) than weekdays (2.26 → 2.23 hours). This trend aligns with earlier observations of lower weekend MAE, likely influenced by COVID-19 schedule shifts and the effect of $\omega_{w}$ on weekend variability. Additional analysis of intra-user variability, its impact, and a cold-start strategy is provided in the supplementary material. Future improvements may incorporate meta-learning or adaptive regularization to better balance global patterns with individual behaviors, and richer contextual signals from large language models may further enhance personalization.

### v2

@@ Introduction
Indeed, the practical efficacy of the DFC algorithm relies on accurate prediction of EV departure time. Inaccurate predictions introduce two critical risks: early charging reduces DFC ’s benefits by prolonging high SOC exposure, whereas late charging can leave insufficient SOC at departure, causing user discomfort (Figure 1). Consequently, departure time emerges as a key control parameter for charging scheduling within the DFC framework. Existing methods primarily rely on historical statistics of usage patterns, offering limited adaptability to dynamic user contexts (frendo2021smart; lindroth2024online). These approaches often overlook real-time causal signals of departure, leading to poor performance and interpretability. For robust deployment of DFC, predictive models must provide temporal precision while adapting to contextual signals that reflect pre-departure behavior.
Pre-departure behavior is shaped by diverse contextual cues, including user behavioral signals such as physical activity, smartphone engagement, and social interactions, as well as environmental factors like ambient light and sound. Smartphone passive sensing provides a scalable and unobtrusive way to capture these signals continuously, enabling a digital phenotyping approach aligned with the quantified-self paradigm (torous2016new). Prior work has shown that machine learning classifiers trained on sliding-window contextual features can improve first-daily departure prediction from home. Combined with DBSCAN, these models outperform historical-pattern baselines (lee2024extending). Yet, framing departure prediction as classification leads to severe label imbalance, which becomes more pronounced at finer time resolutions for real-time inference.
Building on recent advances in deep survival analysis across domains such as healthcare, churn prediction, and mobility, we reformulate departure time prediction as a time-to-event (TTE) problem, applying survival analysis principles to estimate the probability distribution of departure. This reframing offers two key advantages: (1) it mitigates label imbalance by representing each day as a single observation instead of multiple binary windows, and (2) it supports real-time inference where the event time is unknown until departure, enabling probabilistic predictions. Unlike classification, TTE models predict survival functions, providing richer temporal modeling of departure likelihood. To this end, we propose a time-to-departure (TTD) modeling approach based on a Transformer architecture, leveraging its ability to capture temporal dependencies and achieve efficient parallel processing through multi-head attention. The model updates survival probabilities over discretized time grids using real-time contextual signals from passive sensing. This enables timely departure prediction and providing a robust foundation for successful deployment of DFC in real-world environments. Our contributions are threefold:
@@
This discrete-time formulation characterizes the relationship between hazard and survival functions. In our model, we directly estimate the survival probability at each interval, denoted $\hat{S}(t\mid\mathbf{X})$, rather than predicting hazards explicitly.
@@
Our framework predicts $\hat{S}(t\mid\mathbf{X})$ for all intervals in parallel using a Transformer-based architecture, enabling efficient survival estimation with real-time updates. Through self-attention, the model captures both regular daily routine and dynamically changing contextual patterns, supporting timely and adaptive departure predictions (hu2021transformer).
Objective. Given contextual observations up to time $t$, $\mathcal{X}_{0:t}$, the goal is to estimate the sequence of survival probabilities $\{\hat{S}(\tau\mid\mathbf{X})\}_{\tau=0}^{T_{\max}}$, which characterizes the likelihood that departure has not occurred by each interval. Formally, the model produces
$\hat{S}_{0:T_{\max}}=f_{\theta}(\mathcal{X}_{0:t}),$
where $f_{\theta}$ is the Transformer-based architecture parameterized by $\theta$. During training, supervision is applied only up to the observed departure time $t^{*}$, enforcing survival for $t<t^{*}$ and failure at $t=t^{*}$. During inference, the model processes partial sequences in real time, updating survival estimates and triggering departure detection once the predicted survival probability falls below a decision threshold $p$. Figure 2 illustrates this real-time update process.
@@
Output Layer. For each interval $t$, the Transformer produces an output representation $\mathbf{u}_{t}\in\mathbb{R}^{d_{\text{model}}}$, which is passed through a feed-forward network with a sigmoid activation:
$\hat{S}(t\mid\mathbf{X})=\sigma(W_{o}\mathbf{u}_{t}+b_{o}),$
where $\hat{S}(t\mid\mathbf{X})\in(0,1)$ denotes the predicted survival probability $P(T>t\mid\mathbf{X})$. These per-interval estimates form a discrete survival curve that updates in real time as contextual observations evolve throughout the day.
Figure 2: Real-time survival modeling with streaming contextual inputs. During training, the model estimates survival probabilities up to the observed departure event for uncensored sequences. At inference, survival estimates are updated incrementally as new tokens arrive, triggering a departure once the probability falls below a threshold $p$.
@@
We optimize a discrete-time ordinal regression objective inspired by (hu2021transformer). Let the training set contain $N$ user-day sequences indexed by $i\in\{1,\dots,N\}$. For each sequence $i$, let $\mathbf{X}_{i}$ denote the full contextual feature sequence and $t^{*}_{i}\in\{0,\dots,T_{\max}\}$ denote the observed (uncensored) departure time. The model outputs per-interval survival probabilities $\hat{S}(t\mid\mathbf{X}_{i})$ for $t\in[0,T_{\max}]$. Unlike (hu2021transformer), which parameterizes the hazard complement, we directly model $S(t)$ to avoid unnecessary intermediate variables, obtain cleaner uncertainty estimates, and simplify downstream inference for real-time departure detection. This survival estimate forms the basis for the ordinal regression loss used to supervise the model up to the observed event time $t_{i}^{*}$.
Ordinal Regression Loss. Since all training sequences contain an observed departure event, we optimize the negative log-likelihood for uncensored samples. The ordinal regression loss encourages the model to assign high survival probability before the event and a sharp probability drop at the event time:
$\mathcal{L}_{\text{ord},i}=-\sum_{t=0}^{t^{*}_{i}-1}\log\hat{S}(t\mid\mathbf{X}_{i})\;-\;\log\bigl(1-\hat{S}(t^{*}_{i}\mid\mathbf{X}_{i})\bigr).$
The first term encourages the user to remain in the “not departed” state for all intervals before $t^{*}_{i}$, while the second enforces a transition to the departure state at $t^{*}_{i}$. Predictions after the event are excluded, as they are not meaningful in TTD modeling. Since non-departure days represent true non-events rather than right-censored sequences, training is performed solely on samples where a departure is observed.
Soft Supervision with Gaussian Smoothing. Short adjacent intervals (i.e., 5 minutes) may share similar context, making strict supervision sensitive to minor timing shifts near the event. To mitigate this, we apply a normalized Gaussian weighting kernel centered at $t^{*}_{i}$, restricted to pre-event intervals, to smooth the loss around the departure time. This Gaussian-Smoothed Supervision (GSS) enhances robustness to label uncertainty and contextual noise immediately preceding the event:
$w_{i}(t)=\frac{\exp\bigl[-(t-t^{*}_{i})^{2}/(2\sigma^{2})\bigr]}{\sum_{\tau\leq t^{*}_{i}}\exp\bigl[-(\tau-t^{*}_{i})^{2}/(2\sigma^{2})\bigr]},\qquad t\leq t^{*}_{i}.$
The smoothed ordinal regression loss becomes
$\mathcal{L}_{\text{smooth},i}=-\sum_{t=0}^{t^{*}_{i}}w_{i}(t)\Big[$ $\mathbf{1}_{t<t^{*}_{i}}\log\hat{S}(t\mid\mathbf{X}_{i})$
${}+\mathbf{1}_{t=t^{*}_{i}}\log\bigl(1-\hat{S}(t\mid\mathbf{X}_{i})\bigr)\Big].$
Event and Weekend Weights. We incorporate two weighting factors: (1) an event weight $\omega_{e}$, which amplifies the failure term at $t^{*}_{i}$ and scales the corresponding gradients; and (2) a weekend weight $\omega_{w}$, which adjusts sequence-level contributions to account for irregular weekend patterns and stabilize the overall loss.
@@
$\text{where }\ell_{i,t}=\begin{cases}-\log\hat{S}(t\mid\mathbf{X}_{i}),&t<t^{*}_{i},\\[4.0pt] -\omega_{e}\log\bigl(1-\hat{S}(t^{*}_{i}\mid\mathbf{X}_{i})\bigr),&t=t^{*}_{i},\\[4.0pt] 0,&t>t^{*}_{i},\end{cases}$
@@
Figure 3: A diagram of the TTD model architecture. Contextual features and day-of-week embeddings are fused with absolute time features through the alpha-fusion mechanism, with absolute time scaled by a learnable parameter to control its contribution relative to contextual signal. The fused embeddings, combined with positional encoding, are processed by a multi-layer Transformer encoder to capture long-range temporal dependencies. The output layer predicts per-interval survival probabilities $\hat{S}(t\mid\mathbf{X})$ across all intervals in parallel during training and supports real-time inference by updating predictions as new tokens arrive.
@@
Evaluation Metric. Performance is measured using mean absolute error (MAE) in hours, which provides an intuitive measure of temporal deviation. The predicted departure time is taken as the earliest time index at which the estimated survival probability falls below a decision threshold $p$. We vary this $p$ from 0.05 to 0.20 in increments of 0.05 to assess its impact on accuracy, with 0.10 selected as the optimal value through hyperparameter tuning.
@@
Distribution Analysis. Figure 4 presents kernel density estimates (KDE) of predicted versus ground-truth departure times. Historical baselines exhibit high concentration around global averages, particularly in morning and afternoon periods, failing to capture contextual routines relevant to departure. In contrast, our model’s distribution closely tracks the ground-truth curve, demonstrating temporal adaptability despite minor deviations. A slight bias toward earlier predictions is observed, primarily due to the decision threshold $p=0.1$. This suggests that adjusting the $p$ improves practicality of DFC by providing a buffer time before departure.
@@
Figure 5: The impact of hyperparameters: event weight $\omega_{e}$, weekend weight $\omega_{w}$, and survival probability threshold $p$ on the departure time prediction accuracy.
@@
To evaluate the contribution of key design components, we conduct an ablation study along three dimensions: (1) input features, (2) architectural and regularization strategies, and (3) loss design. Table 2 presents MAE (hours) for all days, weekdays, and weekends across these variants.
Feature Contributions. Removing contextual features (w/o context) leads to the largest performance drop (4.47 hours), underscoring the critical role of behavioral and environmental indicators in capturing pre-departure routines. Removing absolute time features (w/o time) results in moderate degradation (3.01 hours), indicating that temporal position cues are important for capturing regular temporal patterns within each day. Excluding day-of-week features (w/o DoW) slightly worsens performance (2.64 hours), highlighting their usefulness in capturing weekly rhythms.
@@
Loss Design and Synergistic Regularization. Removing both time-scale and GSS (w/o t-scale & GSS) increases MAE to 2.36 hours, confirming their synergistic benefit. Time-scale improves temporal calibration across discretized time grids, while GSS provides soft supervision that mitigates over-penalization of near-boundary timing shifts.
@@
We examine the sensitivity of our model to three key hyperparameters: event weight $\omega_{e}$, weekend weight $\omega_{w}$, and the survival probability threshold $p$. Figure 5 presents the MAE heatmaps across different combinations.
Effect of Event Weight. Each subplot corresponds to a fixed value of $\omega_{e}$. Medium values (1.0–1.5) yield the best MAE overall, especially when paired with moderate $\omega_{w}$. Small values weaken the event signal, while large values overemphasize the event point and harm calibration. Notably, the model’s stable performance at $\omega_{e}=1$ suggests that balancing survival and failure signals promotes robust training and stable outcomes. The best results are achieved when $\omega_{e}=1.5$ is combined with $\omega_{w}=1.5$ and a small $p$.
Effect of Weekend Weight. Small $\omega_{w}$ weakens weekend learning, causing the model to fall back on regular weekday patterns and making weekend cues less reliable. With large $\omega_{e}$ and $p$, these weak early cues combined with a softened threshold lead to unstable early predictions. With small $\omega_{e}$ and $p$, the event signal is underemphasized and the strict threshold delays detection, producing late predictions. Thus, small $\omega_{w}$ consistently harms weekend performance, while $\omega_{e}$ and $p$ determine whether errors occur early or late.
Effect of Survival Probability Threshold. $p$ shows a consistent MAE pattern centered around $\omega_{e}=1$, where survival and failure signals are balanced. When $\omega_{e}=1.0\text{–}2.0$, smaller $p$ improves accuracy since the model learns sharper survival drops near the event. When $\omega_{e}=0.5$, smaller $p$ delays detection and worsens performance. In short, strict $p$ works well when the event signal is emphasized.
Interactions and Practical Settings. Overall, moderate $\omega_{e}$ and $\omega_{w}$ (around 1.5) combined with a small $p$ yield the most reliable performance. Small $\omega_{w}$ consistently harms weekend predictions, while extreme $\omega_{e}$ or $p$ shifts the model toward unstable early or late detections. Using moderate weights and a strict threshold avoids these issues and provides accurate, stable real-time departure time predictions.
@@
We further evaluate user-specific adaptation through fine-tuning. All global model parameters remain frozen, and only the last Transformer layer and output layer are updated using user-specific data. Personalization yields a modest overall improvement (MAE: 2.20 → 2.13 hours), with larger gains on weekends (2.07 → 1.85 hours) than weekdays (2.26 → 2.23 hours). This trend aligns with earlier observations of lower weekend MAE, likely influenced by COVID-19 schedule shifts and the effect of $\omega_{w}$ on weekend variability. Additional analysis of intra-user variability, its impact, and a cold-start strategy is provided in the supplementary material. Future improvements may incorporate meta-learning or adaptive regularization to better balance global patterns with individual behaviors, and richer contextual signals from large language models may further enhance personalization.